# HyperBlocker: Accelerating Rule-based Blocking in Entity Resolution using GPUs

### v1

@@ Abstract.
This paper studies rule-based blocking in Entity Resolution (ER). We propose $\mathsf{HyperBlocker}$, a GPU-accelerated system for blocking in ER. As opposed to previous blocking algorithms and parallel blocking solvers, $\mathsf{HyperBlocker}$ employs a pipelined architecture to overlap data transfer and GPU operations. It generates a data-aware and rule-aware execution plan on CPUs, for specifying how rules are evaluated, and develops a number of hardware-aware optimizations to achieve massive parallelism on GPUs. Using real-life datasets, we show that $\mathsf{HyperBlocker}$ is at least 6.8$\times$ and 9.1$\times$ faster than prior CPU-powered distributed systems and GPU-based ER solvers, respectively. Better still, by combining $\mathsf{HyperBlocker}$ with the state-of-the-art ER matcher, we can speed up the overall ER process by at least 30% with comparable accuracy.
@@ 1. Introduction
(a) Execution time
(b) Memory cost on a single GPU
Figure 1. DL-based blocking vs. rule-based blocking
@@
Rule-based blocking in ER has attracted a lot of attention (surveyed in (Li et al., 2020b; Papadakis et al., 2020)). Unfortunately, most of existing rule-based blockers are designed for CPU-based (shared nothing) architectures, leading to unsatisfactory performance. Specifically, a blocker typically conducts pairwise comparisons on all pairs of tuples to obtain candidate matches (Figure 2(a)). In a shared-nothing architecture, data is partitioned and spread across a set of processing units. Each unit independently blocks data using its local memory, which may lead to skewed partitions/computations and rising communication costs, e.g., in Figure 2(c), the first unit is assigned more tuples than the second one and worse still, two tuples that both refer to entity 1 are distributed to different partitions. To avoid missing this match, the second unit has to visit its local memory (Stp. 1) and transfer its data to the first unit (Stp. 2). Then the first unit buffers the data (Stp. 3) and finally, conducts comparison locally (Stp. 4 & 5). The shared memory architecture (Figure 2(b)) is the opposite: all data is accessible from all processing units, allowing for efficient data sharing, collaboration between processing units and dynamic workload scheduling, e.g., the two tuples referred to entity 1 can be directly accessed by both units in Figure 2(b) (Stp. 1 & 2). GPUs are typically based on shared memory architectures, offering promising opportunities to achieve blocking parallelism. However, unlike DL-based blocking approaches, few rule-based methods support the massive parallelism offered by GPUs, despite their greater potential (see Figure 1).
@@ 2. Preliminaries
We first review notations (more in (ful, 2024)) for ER, blocking, and GPUs.
@@
(3) $\varphi_{3}:t.{\mathsf{saddress}}\approx_{\mathsf{ED}}s.{\mathsf{saddress}}\land t.{\mathsf{description}}\approx_{\mathsf{JD}}s.{\mathsf{description}}\allowbreak\to t.{\mathsf{eid}}=s.{\mathsf{eid}}$. It gives another condition for identifying two products, i.e., the two products with similar descriptions are sold from stores with similar addresses are potentially matched.
@@ 3. HyperBlocker: System Overview
Challenges. Existing parallel blocking methods typically rely on multiple CPU-powered machines under the shared nothing architecture, to achieve data partition-based parallelism. They reduce the runtime by using more machines, which, however, is not always feasible due to the increasing communication cost (see Section 1).
@@
Architecture. The architecture of $\mathsf{HyperBlocker}$ is shown in Figure 3. Taken a relation $D$ of tuples and a set $\Delta$ of $\mathsf{MDs}$ discovered offline as input, $\mathsf{HyperBlocker}$ divides the tuples in $D$ into $m$ disjoint partitions and asynchronously processes these data partitions in a pipelined manner, so that the execution at devices and the data transfer can be overlapped. Such architecture helps to mitigate the excessive I/O costs, reduce idle time, and improve hardware utilization.
@@ 4. EPG: Execution Plan Generator
To justify, we tested $\mathsf{PostgreSQL}$ (pos, 2024), a popular DBMS, on different queries, by (a) specifying similarity predicates in different orders and (b) evaluating SQLs with or without OR operators (see details in (ful, 2024)). An in-depth analysis using $\mathsf{PostgreSQL}$’s EXPLAIN feature reveals the following: (a) Although $\mathsf{PostgreSQL}$ may prioritize equality, it does not optimize the order of similarity predicates well, e.g., it evaluates similarity predicates simply as specified in the SQLs, leading to 20% slowdown on average. (b) When handling OR operators, $\mathsf{PostgreSQL}$ may perform sequential scans with nested loops, without utilizing the hash index. These findings highlight the needs of designated optimization for rule-based blocking, where we have to handle similarity comparisons and DNF evaluation efficiently.
@@ 4.2. Execution plan generation
Note that it can be costly to iterate all tuple pairs in $D$ to compute the exact evaluation cost of $p$ on $D$, e.g., on average, it takes more than 100s to compute ${\mathsf{cost}}(p,D)$ on a dataset with 2M tuples (see more in (ful, 2024)). Motivated by this, below we train a shallow NNs, denoted by ${\mathcal{N}}$, (i.e., a small feed-forward neural network (Kraska et al., 2018)) to estimate the exact $T_{p}(t_{1},t_{2})$, since it has been proven effective in approximating a continuous function on a closed interval (Stone, 1937).
Shallow NNs. The inputs of ${\mathcal{N}}$ are two tuples $t_{1}$ and $t_{2}$, and a predicate $p$, that compares the $A$-values of $t_{1}$ and $t_{2}$. It first encodes the attribute type and the $A$-value of $t_{1}$ into an embedding $\vec{t_{1}}$; similarly for $\vec{t_{2}}$. The embeddings are then fed to a feed-forward neural network, which outputs the estimated time for evaluating $p$ on $(t_{1},t_{2})$. We train ${\mathcal{N}}$ offline with training data sampled from historical logs, so that the training data follows the same distribution as $D$ (see (ful, 2024)).
@@
where ${\mathsf{norm}}(\cdot)$ normalizes the estimated cost in the range (0,1].
Remark. As will be verified in Section 6, although $\hat{{\mathsf{cost}}}(p,D)$ is computed by iterating tuple pairs in $D$ (i.e., quadratic cost), it is much faster than computing ${\mathsf{cost}}(p,D)$. Better still, the predicate orderings derived from $\hat{{\mathsf{cost}}}(p,D)$ is close to that derived from ${\mathsf{cost}}(p,D)$.
@@
${\mathsf{sp}}(p,D)={\mathsf{norm}}\big{(}\sqrt{\frac{1}{k}\sum_{i=1}^{k}(b_{i}-\frac{|D|}{k})^{2}}\big{)}.\vspace{-0.7ex}$
@@
${\mathsf{wp}}(\varphi,D)=\prod_{p\in X}{\mathsf{sp}}(p,D),\vspace{-0.7ex}$
@@ 5. Optimizations and Scheduling
While thread divergence is a general issue in GPU-programming, rule-based blocking offers some unique opportunities to mitigate it, e.g., the evaluations of distinct pairs are often independent tasks, making it possible to (a) assign approximately equal tasks to threads, to enable workload balancing, and (b) “steal” tasks from other threads, to cope with different execution paths and data-dependent execution. Below we present the hardware-aware optimization and scheduling techniques that exploit GPU characteristics for massive parallelism, including: (a) efficient device execution of an execution plan (Section 5.1), (b) strategies to mitigate divergence (Section 5.2), (c) collaboration of multiple GPUs (Section 5.3) and more in (ful, 2024).
@@ 5.2. Divergence mitigation strategies
Remark. Note that PSW achieves tuple-level parallelism (i.e., tuples are evaluated in parallel; each thread is responsible for a tuple). There are other design choices for parallelism, e.g., rule-level, predicate-level and token-level parallelisms, so that the entire warp is assigned to a tuple pair and rules, predicates and tokens in text are evaluated in parallel, respectively (one thread per rule/predicate/token, see (ful, 2024) for more). However, these designs may exacerbate divergence, due to, e.g., synchronization cost (e.g., one thread has to wait for results from other threads to make a final decision), or incur unnecessary computation, e.g., when a thread finds a witness $\varphi$ at $(t_{1},t_{2})$ (resp. a predicate $p$ in $\varphi$ such that $h(t_{1},t_{2})\not\models p$), other threads may still check other rules in $\Delta$ (resp. other predicates in $\varphi$).
@@ 5.3. GPU collaboration
In pursuit of this, one can split data evenly so that each GPU handles exactly one (Pandey et al., 2019), or assign multiple partitions to each GPU in a round-robin manner (Rossi and Zhou, 2016). These, however, do not work well since (a) workload can be imbalanced due to skewed execution times of partitions (see statistic in (ful, 2024)), (b) pending partitions may wait when multiple partitions compete for limited PCIe bandwidth or CUDA cores (see Section 2) and (c) they independently conduct blocking on partitions and do not effectively handle scenarios where $t_{i}$ and $t_{j}$ reside on different partitions, resulting in elevated false-negative rates. To address this issue, one can duplicate tuples in multiple partitions (Deng et al., 2022; Chu et al., 2016), but it incurs both memory and data transfer costs.
@@ 6. Experimental Study
We also compared four GPU-based baselines: (4) $\mathsf{DeepBlocker}$ (Thirumuruganathan et al., 2021), (5) $\mathsf{GPUDet}$ (Forchhammer et al., 2013), (6) $\mathsf{Ditto}$ (Li et al., 2020a; cod, 2021b), (7) ${\mathsf{DeepBlocker}}_{\mathsf{Ditto}}$, where $\mathsf{DeepBlocker}$ is the SOTA DL-based blocker, $\mathsf{GPUDet}$ implements well-known similarity algorithms for tuple pair comparison, $\mathsf{Ditto}$ is the SOTA matcher, and ${\mathsf{DeepBlocker}}_{\mathsf{Ditto}}$ uses $\mathsf{DeepBlocker}$ as the blocker and $\mathsf{Ditto}$ as the matcher, respectively. Note that $\mathsf{Ditto}$ takes tuple pairs as input, instead of relations/partitions as other methods. Due to the high cost of $\mathsf{Ditto}$, it is infeasible to feed the Cartesian product of data to $\mathsf{Ditto}$. Thus, for each tuple in ground truths, we adopted a similarity-join method $\mathsf{Faiss}$ (Johnson et al., 2021) to get the top-2 nearest neighbors, as a preprocessing step of $\mathsf{Ditto}$. Denote the resulting baseline by ${\mathsf{Ditto}}_{\mathsf{top2}}$. Since $\mathsf{Faiss}$ often serves as a key component of many ER solutions (Ebraheem et al., 2018; Thirumuruganathan et al., 2021), we also compared $\mathsf{Faiss}$ in (ful, 2024).
@@
Default parameters. Unless stated explicitly, we used the following default parameters, which are best-tuned on each dataset via gird search (Jiménez et al., 2008). Intuitively, gird search is a practical method for systematically exploring the parameter space to get the set of parameters that results in the best performance. The maximum number of predicates in an $\mathsf{MD}$ is 10. The number $m$ of data partitions is given in Table 2. The sizes of intervals and sliding windows, namely $n_{t}$ and $n_{w}$, are 256 and 1024, respectively. For the offline model ${\mathcal{N}}$, we adopted a regression model with 3 hidden layers, with 2, 6, and 1 neurons, respectively. We used ReLU (Nair and Hinton, 2010) as the activation function and Adam (Kingma and Ba, 2015) as the optimizer. We used one GPU by default.
Experimental results. For lack of space, we report our findings on some datasets as follows; consistent on others datasets (more in (ful, 2024)).
@@
(3) Combing $\mathsf{HyperBlocker}$ with ${\mathsf{Ditto}}$, ${\mathsf{HyperBlocker}}_{\mathsf{Ditto}}$ further boosts the accuracy, achieving the best F1-score in $\mathsf{Songs}$. Nevertheless, DL-based solutions still have the best F1-scores in other cases, justifying that none of them can dominate the other in all cases.
@@
Impact of task schedulers. We tested the impact of task schedulers, by comparing $\mathsf{HyperBlocker}$ with two variants, that uses $\mathsf{EvenSplit}$ and $\mathsf{RoundRobin}$ for scheduling (Section 5.3), respectively, by varying $|D|$ in Figure 7(c). $\mathsf{HyperBlocker}$ works better than the two, e.g., when the scale factor is 100%, $\mathsf{HyperBlocker}$ is 1.3$\times$ and 2.2$\times$ faster than $\mathsf{RoundRobin}$ and $\mathsf{EvenSplit}$, respectively, since both variants may limit CUDA’s ability of dynamically scheduling tasks.
@@
(1) Given a noise ratio $\beta\%$, we injected $\beta\%$ noises to training data of ${\mathcal{N}}$, to disturb its distribution, and report RMSE (Root Mean Squared Error), a widely used metric for regression, in Figure 7(f) (the left y-axis). The RMSE of $\mathcal{N}$ does not degrade much when $\beta\%$ = 20%. However, when $\beta\%$ continues to increase, $\mathcal{N}$ becomes inaccurate. A case study about resulting orderings under different $\beta\%$ is given in (ful, 2024).
@@
(3) The average time for computing the actual cost of a predicate is 0.8s on $\mathsf{DBLP}$-$\mathsf{ACM}$, as opposed to 0.007s by ${\mathcal{N}}$ (see more in (ful, 2024)).
@@
Summary. We find the following. (1) $\mathsf{HyperBlocker}$ outperforms prior blockers and integrated ER solutions. It is at least 6.8$\times$, 9.1$\times$, 10.4$\times$, 15.0$\times$, 18.0$\times$, 11.3$\times$ and 15.5$\times$ faster than $\mathsf{SparkER}$, $\mathsf{GPUDet}$, $\mathsf{DeepBlocker}$, $\mathsf{Dedoop}$, $\mathsf{DisDedup}$, $\mathsf{Ditto}$, and ${\mathsf{DeepBlocker}}_{\mathsf{Ditto}}$ respectively. (2) By combining $\mathsf{HyperBlocker}$ with $\mathsf{Ditto}$, we save at least 30% of time with comparable accuracy. (3) $\mathsf{HyperBlocker}$ beats all its variants (except ${\mathsf{HyperBlocker}}_{\mathsf{Ditto}}$) in both runtime and accuracy, justifying the usefulness of various optimizations: (a) EPG specifies an effective evaluation order, improving the runtime by at least 12.4$\times$ and (b) the hardware optimizations on GPUs speedups blocking by at least 3.4$\times$. (3) $\mathsf{HyperBlocker}$ scales well with various parameters, e.g., it completes blocking in 1604s on 36M tuples.

### v2

@@ Abstract.
This paper studies rule-based blocking in Entity Resolution (ER). We propose $\mathsf{HyperBlocker}$, a GPU-accelerated system for blocking in ER. As opposed to previous blocking algorithms and parallel blocking solvers, $\mathsf{HyperBlocker}$ employs a pipelined architecture to overlap data transfer and GPU operations. It generates a data-aware and rule-aware execution plan on CPUs, for specifying how rules are evaluated, and develops a number of hardware-aware optimizations to achieve massive parallelism on GPUs.
Using real-life datasets, we show that $\mathsf{HyperBlocker}$ is at least 6.8$\times$ and 9.1$\times$ faster than prior CPU-powered distributed systems and GPU-based ER solvers, respectively. Better still, by combining $\mathsf{HyperBlocker}$ with the state-of-the-art ER matcher, we can speed up the overall ER process by at least 30% with comparable accuracy.
PVLDB Reference Format:
Xiaoke Zhu, Min Xie, Ting Deng, Qi Zhang. PVLDB, 18(2): 308 - 321, 2024.
doi:10.14778/3705829.3705847 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 2 ISSN 2150-8097.
doi:10.14778/3705829.3705847
@@ 1. Introduction
(a) Execution time
(b) Memory cost
Figure 1. DL-based blocking vs. rule-based blocking
@@
Rule-based blocking in ER has attracted a lot of attention (surveyed in (Li et al., 2020b; Papadakis et al., 2020)). However, most existing rule-based blockers are designed for CPU-based (shared nothing) architectures, leading to unsatisfactory performance. Typically, a blocker conducts pairwise comparisons on all pairs of tuples to obtain candidate matches (Figure 2(a)). In a shared-nothing architecture, data is partitioned and spread across a set of processing units. Each unit independently blocks data using its local memory, which may lead to skewed partitions/computations and rising communication costs, e.g., in Figure 2(c), the first unit is assigned more tuples than the second one and worse still, two tuples that both refer to entity 1 are distributed to different partitions. To avoid missing this match, the second unit has to visit its local memory (Stp. 1) and transfer its data to the first unit (Stp. 2). Then the first unit buffers the data (Stp. 3) and finally, conducts comparison locally (Stp. 4 & 5). The shared memory architecture (Figure 2(b)) is the opposite: all data is accessible from all processing units, allowing for efficient data sharing, collaboration between processing units and dynamic workload scheduling, e.g., the two tuples referred to entity 1 can be directly accessed by both units in Figure 2(b) (Stp. 1 & 2). GPUs are typically based on shared memory architectures, offering promising opportunities to achieve blocking parallelism. However, unlike DL-based approaches, few rule-based methods support the massive parallelism offered by GPUs, despite their greater potential in parallel scalability (see Figure 1).
@@ 2. Preliminaries
We first review the notations for ER, blocking, and the GPU.
@@
(3) $\varphi_{3}:t.{\mathsf{saddress}}\approx_{\mathsf{ED}}s.{\mathsf{saddress}}\land t.{\mathsf{description}}\approx_{\mathsf{JD}}s.{\mathsf{description}}\allowbreak\to t.{\mathsf{eid}}=s.{\mathsf{eid}}$. It gives another condition for identifying two products, i.e., the two products with similar descriptions sold from stores with similar addresses are potentially matched.
@@ 3. HyperBlocker: System Overview
Challenges. Existing parallel blocking methods typically rely on multiple CPU-powered machines under the shared nothing architecture, to achieve data partition-based parallelism. They reduce the runtime by using more machines, which, however, is not always feasible due to, e.g., the increasing communication cost (see Section 1).
@@
Architecture. The architecture of $\mathsf{HyperBlocker}$ is shown in Figure 3. Taken a relation $D$ of tuples and a set $\Delta$ of $\mathsf{MDs}$ discovered offline as input, $\mathsf{HyperBlocker}$ divides the tuples in $D$ into $m$ disjoint partitions and asynchronously processes partitions in a pipelined manner, so that the execution at devices and the data transfer can be overlapped, mitigating the excessive data transfer costs.
@@ 4.2. Execution plan generation
Note that it can be costly to iterate all tuple pairs in $D$ to compute the exact evaluation cost of $p$ on $D$. Thus, below we train a shallow NNs, denoted by ${\mathcal{N}}$, (i.e., a small feed-forward neural network (Kraska et al., 2018)) to estimate the exact $T_{p}(t_{1},t_{2})$, since it has been proven effective in approximating a continuous function on a closed interval (Stone, 1937).
Shallow NNs. The inputs of ${\mathcal{N}}$ are two tuples $t_{1}$ and $t_{2}$, and a predicate $p$, that compares the $A$-values of $t_{1}$ and $t_{2}$. It first encodes the attribute type and the $A$-value of $t_{1}$ into an embedding $\vec{t_{1}}$; similarly for $\vec{t_{2}}$. The embeddings are then fed to a feed-forward neural network, which outputs the estimated time for evaluating $p$ on $(t_{1},t_{2})$. We train ${\mathcal{N}}$ offline with training data sampled from historical logs, so that the training data follows the same distribution as $D$.
@@
where ${\mathsf{norm}}(\cdot)$ normalizes the estimated cost in the range (0,1]. In practice, we can also use a sampled set from $D$ for estimating costs.
@@
${\mathsf{sp}}(p,D)={\mathsf{norm}}\big{(}\sqrt{\frac{1}{k}\sum_{i=1}^{k}(b_{i}-\frac{|D|}{k})^{2}}\big{)}\vspace{-0.7ex}$
Figure 4. Execution tree
@@
${\mathsf{wp}}(\varphi,D)=\prod_{p\in X}{\mathsf{sp}}(p,D)\vspace{-0.7ex}$
@@ 5. Optimizations and Scheduling
While thread divergence is a general issue in GPU-programming, rule-based blocking offers some unique opportunities to mitigate it, e.g., the evaluations of distinct pairs are often independent tasks, making it possible to (a) assign approximately equal tasks to threads, to enable workload balancing, and (b) “steal” tasks from other threads, to cope with different execution paths and data-dependent execution. Below we present the hardware-aware optimization and scheduling techniques that exploit GPU characteristics for massive parallelism, including: (a) efficient device execution of an execution plan (Section 5.1), (b) strategies to mitigate divergence (Section 5.2), (c) collaboration of multiple GPUs (Section 5.3).
@@ 5.3. GPU collaboration
In pursuit of this, one can split data evenly so that each GPU handles exactly one (Pandey et al., 2019), or assign multiple partitions to each GPU in a round-robin manner (Rossi and Zhou, 2016). These, however, do not work well since (a) workload can be imbalanced due to skewed execution times of partitions, (b) pending partitions may wait when multiple partitions compete for limited PCIe bandwidth or CUDA cores (see Section 2) and (c) they independently conduct blocking on partitions and do not effectively handle scenarios where $t_{i}$ and $t_{j}$ reside on different partitions, resulting in elevated false-negative rates. To address these, one can duplicate tuples in multiple partitions (Deng et al., 2022; Chu et al., 2016), but it incurs both memory and data transfer costs.
@@ 6. Experimental Study
We also compared four GPU-based baselines: (4) $\mathsf{DeepBlocker}$ (Thirumuruganathan et al., 2021), (5) $\mathsf{GPUDet}$ (Forchhammer et al., 2013), (6) $\mathsf{Ditto}$ (Li et al., 2020a; cod, 2021b), (7) ${\mathsf{DeepBlocker}}_{\mathsf{Ditto}}$, where $\mathsf{DeepBlocker}$ is the SOTA DL-based blocker, $\mathsf{GPUDet}$ implements well-known similarity algorithms for tuple pair comparison, $\mathsf{Ditto}$ is the SOTA matcher, and ${\mathsf{DeepBlocker}}_{\mathsf{Ditto}}$ uses $\mathsf{DeepBlocker}$ as the blocker and $\mathsf{Ditto}$ as the matcher, respectively. Note that $\mathsf{Ditto}$ takes tuple pairs as input, instead of relations/partitions as other methods. Due to the high cost of $\mathsf{Ditto}$, it is infeasible to feed the Cartesian product of data to $\mathsf{Ditto}$. Thus, for each tuple in GT, we adopted a similarity-join method  (Johnson et al., 2021) to get the top-2 nearest neighbors, as its preprocessing step. Denote the resulting baseline by ${\mathsf{Ditto}}_{\mathsf{top2}}$.
@@
Default parameters. Unless stated explicitly, we used the following parameters, best-tuned on each dataset via gird search (Jiménez et al., 2008). The maximum number of predicates in an $\mathsf{MD}$ is 10. The number $m$ of data partitions is given in Table 2. The sizes of intervals and sliding windows, namely $n_{t}$ and $n_{w}$, are 256 and 1024, respectively. We adopted a regression model as ${\mathcal{N}}$, with 3 hidden layers, with 2, 6, and 1 neurons, respectively. We used ReLU (Nair and Hinton, 2010) as the activation function and Adam (Kingma and Ba, 2015) as the optimizer. We used one GPU by default.
Experimental results. For lack of space, we report our findings on some datasets as follows; consistent on other datasets.
@@
(3) Combing $\mathsf{HyperBlocker}$ with ${\mathsf{Ditto}}$, ${\mathsf{HyperBlocker}}_{\mathsf{Ditto}}$ further boosts the accuracy, achieving the best F1-score in $\mathsf{Songs}$. Nevertheless, DL-based solutions still have the best F1 scores in other cases, justifying that none of them can dominate the other in all cases.
@@
Impact of task schedulers. We tested the impact of task schedulers, by comparing $\mathsf{HyperBlocker}$ with two variants, that uses $\mathsf{EvenSplit}$ and $\mathsf{RoundRobin}$ for scheduling (Section 5.3), respectively, by varying $|D|$ in Figure 7(c). $\mathsf{HyperBlocker}$ works better than the two, e.g., when the scale factor is 100%, $\mathsf{HyperBlocker}$ is 1.3$\times$ and 2.2$\times$ faster than $\mathsf{RoundRobin}$ and $\mathsf{EvenSplit}$, respectively, since both variants may limit CUDA’s ability to dynamically schedule tasks.
@@
(1) Given a noise ratio $\beta\%$, we injected $\beta\%$ noises to training data of ${\mathcal{N}}$, to disturb its distribution, and report RMSE (Root Mean Squared Error), a widely used metric for regression, in Figure 7(f) (the left y-axis). The RMSE of $\mathcal{N}$ does not degrade much when $\beta\%$ = 20%. However, when $\beta\%$ continues to increase, $\mathcal{N}$ becomes inaccurate.
@@
(3) The average time for computing the actual cost of a predicate is 0.8s on $\mathsf{DBLP}$-$\mathsf{ACM}$, as opposed to 0.007s for the estimated cost.
@@
Summary. We find the following. (1) $\mathsf{HyperBlocker}$ outperforms prior blockers and integrated ER solutions. It is at least 6.8$\times$, 9.1$\times$, 10.4$\times$, 15.0$\times$, 18.0$\times$, 11.3$\times$ and 15.5$\times$ faster than $\mathsf{SparkER}$, $\mathsf{GPUDet}$, $\mathsf{DeepBlocker}$, $\mathsf{Dedoop}$, $\mathsf{DisDedup}$, $\mathsf{Ditto}$, and ${\mathsf{DeepBlocker}}_{\mathsf{Ditto}}$ respectively. (2) By combining $\mathsf{HyperBlocker}$ with $\mathsf{Ditto}$, we save at least 30% of time with comparable accuracy. (3) $\mathsf{HyperBlocker}$ beats all its variants (except ${\mathsf{HyperBlocker}}_{\mathsf{Ditto}}$) in both runtime and accuracy, justifying the usefulness of various optimizations: (a) EPG specifies an effective evaluation order, improving the runtime by at least 12.4$\times$ and (b) the hardware optimizations on GPUs speedup blocking by at least 3.4$\times$. (4) $\mathsf{HyperBlocker}$ scales well with various parameters, e.g., it completes blocking in 1604s on 36M tuples.